Circular No. 5800

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Almost 200 precise positions of this comet have now been reported,

about a quarter of them during the past month, notably from CCD images by S. Nakano and by T. Kobayashi in Japan and by E. Meyer, E. Obermair and H. Raab in Austria. These observations are mainly of the “center” of the nuclear train, and this point continues to be the most relevant for orbit computations. Orbit solutions from positions of the brighter individual nuclei will be useful later on, but probably not until the best data can be collected together after the current opposition period. At the end of April, computations by both Nakano and the undersigned were beginning to indicate that the presumed encounter with Jupiter (cf. IAUC 5726, 5744) occurred during the first half of July 1992, and that there will be another close encounter with Jupiter around the end of July 1994. Computations from the May data confirm this conclusion, and the following result was derived by Nakano from 104 observations extending to May 18:

Epoch = 1993 June 22.0 TT

T = 1998 Apr. 5.7514 TT  ω = 22.^{o}9373
e = 0.065832  Ω = 321.5182 }2000.0
q = 4.822184 AU  i = 1.3498
a = 5.162007 AU n^{o} = 0.0840381 P = 11.728 years

This particular computation indicates that the comet’s minimum distance Δ_{J} from the center of Jupiter was 0.0008 AU (i.e., within the Roche limit) on 1992 July 8.8 UT and that Δ_{J} will be only 0.0003 AU (Jupiter’s radius being 0.0005 AU) on 1994 July 25.4.

As noted on IAUC 5726, the positions of the ends of the nuclear train can be satisfied by varying the place in orbit at the time of the 1992 encounter and considering the subsequent differential perturbations. Using the above orbital elements, the undersigned notes that the train as reported on IAUC 5730 corresponds to a variation of ±1.2 seconds. Separation can be regarded as an impulse along the orbit at encounter, although the velocity of separation (or the variation along the orbit) depends strongly on the actual value of Δ_{J}. At the large heliocentric distances involved any differential nongravitational acceleration must be very small, as Z. Sekanina, Jet Propulsion Laboratory, has also noted. Extrapolation to shortly before the 1994 encounter indicates that the train will then be ∼ 20′ long and oriented in p.a. 61^{o}-241^{o}, whereas during the days before encounter the center of the train will be approaching Jupiter from p.a. ∼ 238^{o}.

1993 May 22 Brian G. Marsden